# Foundation Models in Medical Imaging - A Review and Outlook

### v3

@@ 2.1.3 Adapting to Downstream Tasks
Foundation models can generalize across a wide range of downstream tasks, such as image classification, object detection, segmentation, and question answering. Ideally, the model’s latent space is well-structured after pretraining, with the learned embeddings capturing relevant features for various tasks. Methods like PCA and k-nearest neighbors (k-NN) can be used to evaluate how well the embeddings cluster task-specific data, serving as indicators of how well the pretrained model may perform on certain tasks without further adaptation. If the embeddings are well-structured, simple methods like k-NN can yield satisfactory results. However, in most cases, additional adaptation is required to fine-tune the model for specific tasks. This adaptation spectrum ranges from lightweight methods like prompt engineering to more resource-intensive approaches such as full fine-tuning. The choice of adaptation method depends on factors like annotated data availability, computational resources, and the performance requirements for the task.
@@
Prompting can be categorized into zero-shot and few-shot approaches. In zero-shot prompting, the model is given only task instructions without examples. In contrast, few-shot prompting provides a handful of demonstration examples to illustrate the desired behavior before requesting a response. A key technique within few-shot prompting is In-Context Learning (ICL), introduced in GPT-3 [6]. ICL involves providing examples, known as demonstrations, within the input prompt to help the model understand and perform the task at hand. The examples are formatted as input-output pairs, which the model uses to infer the desired behavior and generate appropriate responses for new inputs [32, 33]. Unlike traditional training with explicit labels, ICL helps the model learn from context without retraining, enabling it to generalize to new, unseen tasks. While most ICL methods are in few-shot setting, recently, many-shot ICL was also recently proposed and shown to improve performance [34].
While classic prompt engineering applies mainly to language and vision-language models, there has been emerging research into visual prompting [35, 36]. However, unlike textual prompts, visual prompting techniques often require modifying some model parameters. These approaches typically introduce learnable visual tokens, small trainable modules, or perturbations to the input images that require additional optimization. As such, visual prompting is closer to parameter-efficient fine-tuning than to pure textual prompting, where the model remains entirely unchanged.
@@
If linear probing is insufficient for capturing the complexity of the downstream task, a more powerful adaptation approach involves adding task-specific heads on top of the pretrained model. These smaller neural network modules are fine-tuned for specific tasks, while the foundation model’s weights remain frozen. Task-specific heads can be categorized into those for discriminative tasks, such as image classification and object detection, and generative heads, for tasks like image synthesis or text-to-image generation. For discriminative tasks, these heads may range from simple neural networks like MLPs and CNNs to more complex structures, including additional transformers. Generative heads, on the other hand, may involve autoregressive models, variational autoencoders (VAEs), or diffusion models [37, 38]. This method requires supervised training and labeled data, but the amount of annotated data needed is significantly lower than what would be required to train an entire model from scratch. By focusing fine-tuning efforts on these smaller, task-specific components, computational costs can be reduced.
@@
For scenarios where the pre-trained model needs to be finetuned, but full fine-tuning is too costly, Parameter-Efficient Finetuning (PEFT) techniques aim to reduce computational resources by fine-tuning only a specific set of model parameters [39]. One of the most popular techniques is LoRA (Low-Rank Adaptation), which focuses on updating a set of low-rank matrices [40, 41, 42]. Another widely used PEFT method is adapter tuning, where small adapter modules are inserted into the layers of a pretrained model and trained for specific tasks while keeping the original model parameters frozen [43]. This allows for efficient task-specific adaptation without requiring full model updates. Adapter tuning is particularly useful in multitask learning, where a single model is adapted to multiple tasks without retraining the entire network. Another approach is prompt tuning, where a small set of trainable prompt embeddings is optimized while keeping the model’s weights frozen [44]. Unlike traditional prompt engineering, which involves manually crafting text-based prompts, prompt tuning learns task-specific soft prompts in a data-driven manner, making it a more scalable adaptation method.
@@
Full fine-tuning involves updating all of the model’s parameters based on the labeled data for a specific downstream task. For example, in a classification task, the model would not only adjust the final classification layer but also modify its internal representations to better capture the features specific to that task. This approach typically leads to the best performance for the task at hand but comes with higher computational costs, as all parameters are updated during training. Additionally, sufficient labeled data is required to prevent overfitting and ensure that the model generalizes well to new examples [45].
Finally, some specialized techniques that involve full fine-tuning exist. For instance, Instruction-Tuning uses demonstration-based instructions during the model’s training phase, updating its parameters to optimize it for better task-following at inference time [46]. A vision-language extension of this approach is Visual Instruction Tuning, as implemented in LLaVA [47], where vision-language models are fine-tuned using multimodal instruction-response pairs to improve their ability to follow natural language prompts grounded in images.
Another notable method is Reinforcement Learning from Human Feedback (RLHF) [48]. RLHF involves fine-tuning models based on feedback from human users, which can be particularly useful for aligning model outputs with human preferences and improving performance on subjective tasks.
@@ 3 Pathology
Pathology involves distinguishing diseases based on clinical, macroscopic, microscopic, and molecular characteristics by comparing them with predefined criteria. A pathology report provides a diagnosis but also includes information on prognostic (disease behavior) and predictive factors (response to therapy) [49]. Computational pathology has incorporated many downstream tasks from this workflow and continues to develop improved solutions for them using different methods. In computational pathology, whole slide images (WSIs) are widely used to train AI models to improve cancer diagnosis accuracy, discover new biomarkers, and tailor treatment strategies according to individual patient profiles and specific tumor characteristics [50]. WSIs are obtained by digitizing conventional pathology glass slides using slide scanners, which contain thin tissue sections typically stained with hematoxylin and eosin (H&E) or immunohistochemistry (IHC) stains. During this scanning process, the entire slide is captured at different magnification levels, providing a comprehensive view of the tissue sample, from a broad overview to the fine details of individual cells. The gigapixel size of the resulting WSIs makes it difficult for current hardware to process them entirely due to computation and memory constraints. Therefore, WSIs, at a fixed Microns Per Pixel (MPP), are usually split into smaller tiled patches, which are fed to the model individually.
Several factors can influence image analysis on WSIs. One major factor is the manual processing during preanalytical steps in a pathology laboratory, as it can introduce artifacts due to slight procedural differences [51]. Additionally, variations in H&E staining protocols across laboratories and using different slide scanners contribute to color inconsistencies between WSIs from different centers [52], highlighting the need for robust models to overcome these variances.
@@ 3.1 Towards Pathology Foundation Models
In digital pathology image analysis, deep learning methods typically follow a two-stage pipeline: a pre-trained backbone encoder computes lower-dimensional embeddings for individual tiles, which are then processed by a task-specific head. Segmentation tasks are an exception, where nnU-Net models are commonly used at the patch level. Encoder weights are often kept frozen to avoid overfitting on small labeled datasets, as the target labels are usually slide-level, such as clinical outcomes or molecular profiles. In practice, freezing also helps reduce computational demands, since end-to-end training from patch to slide level can be memory-intensive. These tasks are typically framed as weakly supervised problems, with multi-instance learning used to aggregate tile-level representations.
@@
Initially, transfer learning was employed to initialize encoders with weights from natural image datasets like ImageNet [53]. However, recent efforts have shifted toward SSL in the pathology domain, unlocking the potential for in-domain pretraining and significantly improving performance. Some studies applied contrastive SSL frameworks [54, 55, 56], while others explored pathology-oriented pretext tasks based on WSI characteristics like magnification or cross-stain prediction [57, 58]. Early pre-training often focused on tissue types similar to downstream tasks, but more diverse pan-cancer datasets are now being used to train general-purpose encoders. These encoders serve as building blocks for various pathology tasks, with both public datasets like TCGA [59] and proprietary data being used in this trend of scaling up data and model size.
@@ 3.2 Pathology Foundation Models
This section reviews recent advances in FMs for pathology, focusing on tile-level foundation models, which follow the two-stage pipeline outlined in Section 3.1. We first discuss the development of tile-level models in chronological order, grouped by SSL frameworks. Next, we cover vision-language FMs, slide-level models, and the application of SAM for pathology data. An overview of landmark papers is provided in Tables 1, 2 and 3.
@@ 3.2.1 Contrastive Learning
[60] were the first to utilize a large number and diverse set of histopathology images for self-supervised training. Public data from TCGA, CPTAC111 https://cptac-data-portal.georgetown.edu/cptacPublic, and several public challenge datasets were merged into a unified pretraining set with 25 thousand full WSIs and more than 206 thousand image patches, including different organs, staining types, and resolutions. Downstream performance on several classification, regression, and segmentation tasks revealed that in-domain pretraining of CNNs with SimCLR on a larger and more diverse dataset could improve the quality of the learned embeddings.
[61] used large-scale supervised transfer learning on natural images to initialize the backbone encoder before finetuning it on domain-specific unlabeled data using SimCLR. This training strategy, dubbed REMEDIS, achieved better performance on tasks such as metastasis detection and survival prediction and more data-efficient generalization to out-of-distribution data compared to supervised baselines pre-trained on natural images only.
@@
A challenge in contrastive learning is defining informative contrastive pairs. Typically, two augmented views of the same instance are considered a positive pair, and those from different instances are negative pairs. In histopathology, where patches of WSIs are often semantically similar, this approach can be problematic. To address this, [62] proposed cluster-guided contrastive learning (CCL) to reduce false-negative samples. [63] introduced semantically relevant contrastive learning (SRCL), an SSL strategy based on MocoV3 that uses latent space similarity to identify additional positive pairs. Using SRCL, a hybrid backbone integrating a CNN and multi-scale Swin Transformer architecture, coined CTransPath, was trained on a large set of 32 thousand public pan-cancer WSIs from TCGA and PAIP [^2] http://www.wisepaip.org/paip. CTransPath was one of the first transformer-based feature extractors optimized using SSL at such a scale.
@@ 3.2.2 Self-distillation with DINO
Several studies later adopted the DINO self-distillation framework for ViT-based pathology FMs. [64] introduced the Hierarchical Image Pyramid Transformer (HIPT), which models the hierarchical structure of WSIs by aggregating visual tokens across different scales using DINO. This enables the capture of coarser-grained features across a larger context. [65] applied DINO to a proprietary dataset with 3 billion patches from 432,000 WSIs, substantially larger than those in previous studies. [66] benchmarked four representative SSL methods, including DINO, and found that in-domain pre-training consistently outperformed supervised pre-training with ImageNet, though no SSL method showed a significant advantage. [67] used DINO to train HistoEncoder, an FM for prostate cancer that utilizes the XCiT backbone [68], reducing complexity compared to traditional ViTs. HistoEncoder outperformed models pre-trained on natural images for cancer classification and mortality prediction, even without fine-tuning or with less data, making it more accessible to users with limited resources.
@@ 3.2.3 Masked Image Modeling
Inspired by the success of MIM in the natural image domain, most recent studies on pathology FMs use this SSL paradigm for in-domain pre-training. While [69] introduced a BEiT-based pathology FM, the iBOT and DINOv2 frameworks are most commonly employed by other works that we discuss in more detail below.
@@
[70] analyzed the scalability of iBOT and the impact of data diversity by pretraining different ViT model sizes on either colon adenocarcinoma data or (a selection of) pan-cancer data from TCGA, which eventually resulted in the FM known as Phikon. A large panel of 17 slide-level and tile-level downstream tasks, including the prediction of histological and molecular subtypes, genomic alterations, and survival, was used for evaluation. Phikon improved performance over previously published methods on all patch-level tasks and most slide-level classification tasks with, on the latter, an average performance increase of +1.4% and +4.0% compared to CTransPath and HIPT, respectively.
Table 1: Tile-Level Vision Encoder Pathology FMs. The data size for pre-training is given in number of patches and the number of WSIs in parentheses.
@@
  • a
    Proprietary data
  • b
    Available upon request
  • c
    Partly available
@@
Around the same time, [71] introduced UNI, a ViT-Large model trained with DINOv2 on the Mass-100k dataset, which included over 100,000 WSIs across 20 pathology types. UNI was compared with CTransPath and REMEDIS on a range of tasks, including slide-level classification and ROI-level tasks (classification, segmentation, retrieval, and prototyping) across 34 anatomic pathology tasks. UNI outperformed other models on average but performed similarly to REMEDIS in ROI segmentation. UNI also demonstrated better label efficiency and resolution-agnosticity in classification tasks.
@@
Shortly after, [72] published Virchow, a ViT-Huge model trained on a proprietary dataset of 2 billion patches from 1.5 million WSIs and 119,000 patients, making it the largest pathology FM at the time. Virchow outperformed UNI, Phikon, and CTransPath on public benchmark tile-level tasks and pan-cancer detection. Though it performed better in biomarker prediction, the differences were not significant. Notably, in detecting prostate cancer, invasive breast cancer, and breast cancer metastasis to lymph nodes, Virchow was competitive with specialist commercial models, despite using a less refined training set.
@@
Building on [72], [73] introduced domain-inspired augmentations and regularization to improve Virchow’s training. The dataset was expanded to 3.1 million WSIs from 225,000 patients across various institutions and tissue types. The resulting Virchow2 model improved performance on in-distribution and out-of-distribution benchmarks, and further improvements were seen with the larger Virchow2G model, using a ViT-G architecture. More recently, building on Phikon, Phikon-v2 was proposed, which was trained on 58.4K publically available WSIs with DinoV2 [74]. Phikon-v2 outperformed Phikon, and its performance was on-par with that of FMs trained on large proprietary datasets.
@@
Besides the aforementioned studies, several other pathology FMs have been pre-trained with DINOv2, such as RudolfV [75], PLUTO [76], Hibou-B/Hibou-L [77], and H-optimus-O [78]. Finally, Atlas [79] is a ViT-H/14 pathology foundation model trained using a DINOv2-style self-supervised learning framework adapted from RudolfV. Unlike prior models that typically rely on H&E-stained slides at a single magnification, Atlas was trained on a highly diverse dataset of 1.2 million WSIs from 490,000 cases, containing 70 tissue types, multiple magnifications, scanner types, and staining protocols, including H&E, IHC, and special stains. Despite having a moderate size of 632 million parameters, which is the same as Virchow2 and significantly fewer than Prov-GigaPath’s 1.1 billion, Atlas achieves the highest average performance across 21 public benchmarks as well as additional tumor-microenvironment and cancer typing tasks. It outperforms both Virchow2 and H-Optimus-0 by 1.1 percentage points, suggesting that broad data diversity combined with strong self-supervised learning mightbe favored over simply scaling in both model and dataset size (see also 6.1.3).

### v4

@@ 2.1.3 Adapting to Downstream Tasks
Foundation models can generalize across a wide range of downstream tasks, such as image classification, object detection, segmentation, and question answering. Ideally, the model’s latent space is well-structured after pretraining, with the learned embeddings capturing relevant features for various tasks. Methods like PCA and k-nearest neighbors (k-NN) can be used to evaluate how well the embeddings cluster task-specific data, serving as indicators of how well the pretrained model may perform on certain tasks without further adaptation. If the embeddings are well-structured, simple methods like k-NN can yield satisfactory results. However, in most cases, additional adaptation is required to fine-tune the model for specific tasks. Adaptation methods can be viewed along a spectrum of increasing resource requirements: from approaches that leave model parameters untouched (prompting, linear probing), to those that add lightweight task-specific components while keeping the foundation model frozen (task-specific heads, parameter-efficient fine-tuning), and finally to approaches that update all parameters (full fine-tuning). The choice of adaptation method depends on factors like annotated data availability, computational resources, and the performance requirements for the task.
@@
Prompting can be categorized into zero-shot and few-shot approaches. In zero-shot prompting, the model is given only task instructions without examples. In contrast, few-shot prompting provides a handful of demonstration examples to illustrate the desired behavior before requesting a response. A key technique within few-shot prompting is In-Context Learning (ICL), introduced in GPT-3 [brown_language_2020]. ICL involves providing examples, known as demonstrations, within the input prompt to help the model understand and perform the task at hand. The examples are formatted as input-output pairs, which the model uses to infer the desired behavior and generate appropriate responses for new inputs [dong2022survey, min2022rethinking]. Unlike traditional training with explicit labels, ICL helps the model learn from context without retraining, enabling it to generalize to new, unseen tasks. While most ICL methods are in few-shot setting, recently, many-shot ICL was also proposed and shown to improve performance [agarwal2024many].
While classic prompt engineering applies mainly to language and vision-language models, there has been emerging research into visual prompting [jia2022visual, bahng2022exploring]. However, unlike textual prompts, visual prompting techniques often require modifying some model parameters. These approaches typically introduce learnable visual tokens, small trainable modules, or perturbations to the input images that require additional optimization. As such, visual prompting is closer to parameter-efficient fine-tuning than to pure textual prompting, where the model remains entirely unchanged.
@@
If linear probing is insufficient for capturing the complexity of the downstream task, a more powerful adaptation approach involves adding task-specific heads on top of the pretrained model. These smaller neural network modules are fine-tuned for specific tasks, while the foundation model’s weights remain frozen. Task-specific heads can be categorized into those for discriminative tasks, such as image classification and object detection, and generative heads, for tasks like image synthesis or text-to-image generation. For discriminative tasks, these heads may range from simple neural networks like MLPs and CNNs to more complex structures, including additional transformers. Generative heads, on the other hand, may involve autoregressive models, variational autoencoders (VAEs), or diffusion models [kingma2022autoencoding, ho2020denoising]. These methods require supervised training and labeled data, but the amount of annotated data needed is significantly lower than what would be required to train an entire model from scratch. By focusing fine-tuning efforts on these smaller, task-specific components, computational costs can be reduced.
@@
For scenarios where the pre-trained model needs to be finetuned, but full fine-tuning is too costly, Parameter-Efficient Finetuning (PEFT) techniques aim to reduce computational resources by fine-tuning only a specific set of model parameters [tay2023efficient]. One of the most popular techniques is LoRA (Low-Rank Adaptation), which focuses on updating a set of low-rank matrices [hu2021lora, zhu2023melo, lian2024less]. Another widely used PEFT method is adapter tuning, where small adapter modules are inserted into the layers of a pretrained model and trained for specific tasks while keeping the original model parameters frozen [houlsby2019parameter]. This allows for efficient task-specific adaptation without requiring full model updates. Adapter tuning is particularly useful in multitask learning, where a single model is adapted to multiple tasks without retraining the entire network. Another approach is prompt tuning, where a small set of trainable prompt embeddings is optimized while keeping the model’s weights frozen [lester2021power]. Unlike traditional prompt engineering, which involves manually crafting text-based prompts, prompt tuning learns task-specific soft prompts in a data-driven manner, making it a more scalable adaptation method.
@@
Full fine-tuning involves updating all of the model’s parameters based on the labeled data for a specific downstream task, allowing the model to adjust both its internal representations and the final task-specific layers. This approach typically leads to the best performance for the task at hand but comes with higher computational costs. Additionally, sufficient labeled data is required to prevent overfitting and ensure that the model generalizes well to new examples [howard2018universal].
Finally, some specialized techniques that involve full fine-tuning have been developed to improve task-following and alignment with human preferences. For instance, instruction-tuning uses demonstration-based instructions during the model’s training phase, updating its parameters to optimize it for better task-following at inference time [wei2021finetuned]. A vision-language extension of this approach is visual instruction tuning, as implemented in LLaVA [liu2023visual], where vision-language models are fine-tuned using multimodal instruction-response pairs to improve their ability to follow natural language prompts grounded in images. Furthermore, Reinforcement Learning from Human Feedback (RLHF) [kaufmann2023survey] incorporates human evaluations by training a reward model on preference data, which then guides reinforcement learning to align model outputs with human preferences. This enables more nuanced, context-dependent behaviors that would have been difficult to capture with standard supervised objectives. Both instruction tuning and RLHF are also central to developing more agentic capabilities in models: meaning, the capacity for autonomous planning and the execution of multi-step tasks [acharya2025agentic]. These techniques provide the necessary alignment with demonstrations and human preferences to support such goal-directed behavior. When combined with chain-of-thought prompting [wei2022chain], where the model explicitly generates intermediate reasoning steps, agentic FMs can perform complex multi-step reasoning, plan sequences of actions, and maintain context across tasks.
@@ 3 Pathology
Pathology involves distinguishing diseases based on clinical, macroscopic, microscopic, and molecular characteristics by comparing them with predefined criteria. A pathology report provides a diagnosis but also includes information on prognostic (disease behavior) and predictive factors (response to therapy) [funkhouser2020pathology]. Computational pathology has incorporated many downstream tasks from this workflow and continues to develop improved solutions for them using different methods. Central to these tasks are whole slide images (WSIs), which are widely used to train AI models to improve cancer diagnosis accuracy, discover new biomarkers, and tailor treatment strategies according to individual patient profiles and specific tumor characteristics [song2023artificial]. WSIs are obtained by digitizing conventional pathology glass slides using slide scanners, which contain thin tissue sections typically stained with hematoxylin and eosin (H&E) or immunohistochemistry (IHC) stains. During this scanning process, the entire slide is captured at different magnification levels, providing a comprehensive view of the tissue sample, from a broad overview to the fine details of individual cells. The gigapixel size of the resulting WSIs makes it difficult for current hardware to process them entirely due to computation and memory constraints. Therefore, WSIs, at a fixed Microns Per Pixel (MPP), are usually split into smaller tiled patches, which are fed to the model individually.
Several factors can influence image analysis on WSIs. One major factor is the manual processing during preanalytical steps in a pathology laboratory, as it can introduce artifacts due to slight procedural differences [taqi2018review]. Additionally, variations in H&E staining protocols across laboratories and using different slide scanners contribute to color inconsistencies between WSIs from different centers [ochi2024registered], highlighting the need for robust models to overcome these variances.
@@ 3.1 Towards Pathology Foundation Models
In digital pathology image analysis, deep learning methods typically follow a two-stage pipeline: a pre-trained backbone encoder computes lower-dimensional embeddings for individual tiles, which are then processed by a task-specific head. Segmentation tasks are an exception: nnU-Net models are commonly applied at the patch level, and encoder weights are often kept frozen to prevent overfitting when patch-level labels are limited. In contrast, slide-level prediction tasks, such as predicting clinical outcomes or molecular profiles, typically involve large sets of patches with only one label per slide. These tasks are framed as weakly supervised problems, where multiple-instance learning (MIL) is used to aggregate patch-level embeddings into a slide-level prediction. Freezing encoder weights in this context also helps reduce computational demands, since end-to-end training from patches to slide-level predictions can be memory-intensive.
@@
Initially, transfer learning was employed to initialize encoders with weights from natural image datasets like ImageNet [saillard2020predicting]. However, recent efforts have shifted toward SSL in the pathology domain, unlocking the potential for in-domain pretraining and significantly improving performance. Some studies applied contrastive SSL frameworks [lu2019semi, dehaene2020self, zhao2020predicting], while others explored pathology-oriented pretext tasks based on WSI characteristics like magnification or cross-stain prediction [koohbanani2021self, yang2021self]. Early pre-training often focused on tissue types similar to downstream tasks, but more diverse pan-cancer datasets are now being used to train general-purpose encoders. These encoders serve as building blocks for various pathology tasks, with both public datasets like TCGA [weinstein2013cancer] and proprietary data being used in this trend of scaling up data and model size.
@@ 3.2 Pathology Foundation Models
This section reviews recent advances in FMs for pathology, focusing on tile-level foundation models, which follow the two-stage pipeline outlined in Section 3.1. We first discuss the development of tile-level models, grouped by SSL frameworks. Next, we cover vision-language FMs, slide-level FMs, and the application of SAM for pathology data. An overview of landmark papers is provided in Tables 1, 2 and 3.
@@ 3.2.1 Contrastive Learning
[ciga2022self] were the first to utilize a large number and diverse set of histopathology images for self-supervised training. Public data from TCGA, CPTAC [^1] https://cptac-data-portal.georgetown.edu/cptacPublic, and several public challenge datasets were merged into a unified pretraining set with 25 thousand full WSIs and more than 206 thousand image patches, including different organs, staining types, and resolutions. Downstream performance on several classification, regression, and segmentation tasks revealed that in-domain pretraining of CNNs with SimCLR on a larger and more diverse dataset could improve the quality of the learned embeddings.
[azizi_robust_2023] used large-scale supervised transfer learning on natural images to initialize the backbone encoder before finetuning it on domain-specific unlabeled data using SimCLR. This training strategy, dubbed REMEDIS, achieved better performance on tasks such as metastasis detection and survival prediction and more data-efficient generalization to out-of-distribution data compared to supervised baselines pre-trained on natural images only.
@@
A challenge in contrastive learning is defining informative contrastive pairs. Typically, two augmented views of the same instance are considered a positive pair, and those from different instances are negative pairs. In histopathology, where patches of WSIs are often semantically similar, this approach can be problematic. To address this, [wang2023retccl] proposed cluster-guided contrastive learning (CCL) to reduce false-negative samples. [wang2022transformer] introduced semantically relevant contrastive learning (SRCL), an SSL strategy based on MocoV3 that uses latent space similarity to identify additional positive pairs. Using SRCL, a hybrid backbone integrating a CNN and multi-scale Swin Transformer architecture, coined CTransPath, was trained on a large set of 32 thousand public pan-cancer WSIs from TCGA and PAIP [^2] http://www.wisepaip.org/paip. CTransPath was one of the first transformer-based feature extractors optimized using SSL at such a scale.
@@ 3.2.2 Self-distillation with DINO
Several studies later adopted the DINO self-distillation framework for ViT-based pathology FMs. [chen2022scaling] introduced the Hierarchical Image Pyramid Transformer (HIPT), which models the hierarchical structure of WSIs by aggregating visual tokens across different scales using DINO. This enables the capture of coarser-grained features across a larger context. [campanella2023computational] applied DINO to a proprietary dataset with 3 billion patches from 432,000 WSIs, substantially larger than those in previous studies. [kang2023benchmarking] benchmarked four representative SSL methods, including DINO, and found that in-domain pre-training consistently outperformed supervised pre-training with ImageNet, though no SSL method showed a significant advantage. [pohjonen2024histoencoderdigitalpathologyfoundation] used DINO to train HistoEncoder, an FM for prostate cancer that utilizes the XCiT backbone [elnouby2021xcitcrosscovarianceimagetransformers], reducing complexity compared to traditional ViTs. HistoEncoder outperformed models pre-trained on natural images for cancer classification and mortality prediction, even without fine-tuning or with less data, making it more accessible to users with limited resources.
@@ 3.2.3 Masked Image Modeling
Inspired by the success of MIM in the natural image domain, most recent studies on pathology FMs use this SSL paradigm for in-domain pre-training. While [yang2024foundation] introduced a BEiT-based pathology FM, the iBOT and DINOv2 frameworks are most commonly employed by other works that we discuss in more detail below.
@@
[filiot2023scaling] analyzed the scalability of iBOT and the impact of data diversity by pretraining different ViT model sizes on either colon adenocarcinoma data or (a selection of) pan-cancer data from TCGA, which eventually resulted in the FM known as Phikon. A large panel of 17 slide-level and tile-level downstream tasks, including the prediction of histological and molecular subtypes, genomic alterations, and survival, was used for evaluation. Phikon improved performance over previously published methods on all patch-level tasks and most slide-level classification tasks, with average gains of +1.4% over CTransPath and +4.0% over HIPT. These improvements were partly attributed to iBOT’s masked image modeling SSL method, which enabled richer representations than the contrastive (CTransPath) and distillation-based (HIPT) approaches.
Table 1: Tile-Level Vision Encoder Pathology FMs. Backbone and SSL objective show how the model was pretrained. The data size for pre-training is given in number of patches and the number of WSIs in parentheses. Availability shows whether the weights and pretraining data are available online.
@@
a Proprietary data  b Available upon request  c Partly available
@@
Around the same time, [chen2024towards] introduced UNI, a ViT-Large model trained with DINOv2 on the Mass-100k dataset, which included over 100,000 WSIs across 20 pathology types. UNI was compared with CTransPath and REMEDIS on a range of tasks, including slide-level classification and ROI-level tasks (classification, segmentation, retrieval, and prototyping) across 34 anatomic pathology tasks. UNI outperformed other models on average but performed similarly to REMEDIS in ROI segmentation. UNI also demonstrated better label efficiency and resolution-agnosticity in classification tasks.
@@
Shortly after, [vorontsov2024foundation] published Virchow, a ViT-Huge model trained on a proprietary dataset of 2 billion patches from 1.5 million WSIs and 119,000 patients, making it the largest pathology FM at the time. Virchow outperformed UNI, Phikon, and CTransPath on public benchmark tile-level tasks and pan-cancer detection. Though it performed better in biomarker prediction, the differences were not significant. Notably, in detecting prostate cancer, invasive breast cancer, and breast cancer metastasis to lymph nodes, Virchow was competitive with specialist commercial models, despite using a less refined training set.
@@
Building on [vorontsov2024foundation], [zimmermann2024virchow] introduced domain-inspired augmentations and regularization to improve Virchow’s training. The dataset was expanded to 3.1 million WSIs from 225,000 patients across various institutions and tissue types. The resulting Virchow2 model improved performance on in-distribution and out-of-distribution benchmarks, and further improvements were seen with the larger Virchow2G model, using a ViT-G architecture. More recently, building on Phikon, Phikon-v2 was proposed, which was trained on 58.4K publically available WSIs with DinoV2 [filiot2024phikon]. Phikon-v2 outperformed Phikon, and its performance was on-par with that of FMs trained on large proprietary datasets.
@@
Besides the aforementioned studies, several other pathology FMs have been pre-trained with DINOv2, such as RudolfV [dippel2024rudolfv], PLUTO [juyal2024pluto], Hibou-B/Hibou-L [nechaev2024hibou], and H-optimus-O [hoptimus0]. Finally, Atlas [alber2025novel] is a ViT-H/14 pathology foundation model trained using a DINOv2-style self-supervised learning framework adapted from RudolfV. Unlike prior models that typically rely on H&E-stained slides at a single magnification, Atlas was trained on a highly diverse dataset of 1.2 million WSIs from 490,000 cases, containing 70 tissue types, multiple magnifications, scanner types, and staining protocols, including H&E, IHC, and special stains. Despite having a moderate size of 632 million parameters, which is the same as Virchow2 and significantly fewer than Prov-GigaPath’s 1.1 billion, Atlas achieves the highest average performance across 21 public benchmarks as well as additional tumor-microenvironment and cancer typing tasks. It outperforms both Virchow2 and H-Optimus-0 by 1.1%, suggesting that broad data diversity combined with strong self-supervised learning might be favored over simply scaling in both model and dataset size (see also 6.1.3).